John invested the following amounts in three stocks:

Security Investment Beta
Stock A $488,268 2.45
Stock B $796,341 1.87
Stock C $468,348 2.18

Calculate the beta portfolio.
Answer #1

Total of investment = 488268 + 796341 + 468348 = 1752957

Now,

Beta of portfolio = Stock A * Stock A beta / Total investment + Stock B * Stock B beta / Total investment + Stock C * Stock C beta / Total investment

Beta of portfolio = 488268 * 2.45 / 1752957 + 796341* 1.87 / 1752957 + 468,348*2.18 / 1752957

Beta of portfolio = 0.682422102 + 0.849511808 + 0.582443631

Beta of the portfolio = 2.11437754

Beta of the portfolio = 2.11